if the endpoints of a line segment are (14, -8) and (4,12) what would the midpoint be? use the midpoint formula: (Xm, Ym)= (x1+x2, y1+y2)

2 2


1
Expert's answer
2021-09-30T17:47:40-0400

Given the endpoints of a line segment are (14,-8) and (4,12)

This means "x_1=14,y_1=-8,x_2=4,y_2=12"

Mid point formula:(cm,my)="(\frac{x_1+x_2}{2}),(\frac{y_1+y_2}{2})=
(\frac{14+4}{2}),(\frac{-8+12}{2})\\
=(\frac{16}{2}),(\frac{4}{2})\\=(8,2)"

The midpoint of the line segment will be (8,2)